OpenStudy (anonymous):

2 radical 45 + 5 radical 125

OpenStudy (tkhunny):

45 = 5*9 125 = 5*25

OpenStudy (anonymous):

yea?

OpenStudy (tkhunny):

That is your hint. Use it.

OpenStudy (anonymous):

i know that but i never learned radicals

OpenStudy (anonymous):

break up radical 45 and radical 125

OpenStudy (anonymous):

hw

OpenStudy (tkhunny):

Then why were you given this problem? Find perfect squares and pull them outside the radical expression. \(\sqrt{25} = 5\) \(\sqrt{5\cdot25} = \sqrt{5}\cdot\sqrt{25} = \sqrt{5}\cdot5\)

OpenStudy (anonymous):

im studying for my final but i never made it to that chapter

OpenStudy (tkhunny):

Then you're probably toast. How can anyone help you with no background whatsoever?

OpenStudy (anonymous):

can u just give me the steps so i can understand?

OpenStudy (tkhunny):

Pretty much did. Sorry.